OpenStudy (anonymous):

I need help with inverse matrix, plz let matrix A= -2 -7 -9 2 5 6 1 3 4 find the second column of inverse matrix A^-1 without computing the other columns.

OpenStudy (amistre64):

cofactor, and divide by determinant

OpenStudy (amistre64):

and a transpose ... so its really the setup of row 2 cofactors, divided by |A|, transposed

OpenStudy (amistre64):

-2 -7 -9 --------- a b c --------- 1 3 4 a = - |-7 -9| | 3 4| -------- |A| b = |-2 -9| | 1 4| -------- |A| c = - |-2 -7| | 1 3| -------- |A| [a b c] is the column vector of A^-1

OpenStudy (amistre64):

- (-45-36-56) -2 -7 -9 -2 -7 2 5 6 2 5 1 3 4 1 3 -40-42-54 which of course is eqaul to 1 for the determinant ...

OpenStudy (anonymous):

so the 2nd column of the inverse matrix is related to 2nd row of the A matrix?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

in inverse can be formulated from taking the cofactors of the given matrix, dividing by the determinant, and transposing

OpenStudy (amistre64):

|a b| |c d| a to |d| = d b to -|c| = -c c to -|b| = -b d to |a| = a to give us a cofactor matrix of: | d -c| |-b a| divide the elements by the determinant; lets assume 1 for simplicity here; and transpose | d -b| |-c a| this would be the inverse matrix of a 2x2

OpenStudy (amistre64):

since the determinant of yours is graciously 1 ... this simplifies to a = - |-7 -9| = 1 | 3 4| b = |-2 -9| = 1 | 1 4| c = - |-2 -7| = -1 | 1 3|

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

does it make sense?

OpenStudy (anonymous):

sure